OpenStudy (ulahlynn):

A parabola has directrix y=-3, axis x=4 and latus rectum equal to 2. Find its equation. pls help :)

OpenStudy (amistre64):

i can yes ... do you recall what a latus rectum is?

OpenStudy (ulahlynn):

uhmm. its a chord that passes through the focus? hahaha.. im not that sure though. :D

OpenStudy (amistre64):

youre right actually :) i have an idea on how to do this

OpenStudy (amistre64):

lets create a parabola that fits, then move it into place

OpenStudy (ulahlynn):

ok. :)

OpenStudy (amistre64):

this distance between focus and directrix should be 6 ... so lets have a parabola of vertex 0,-3 with roots of -1,0 and 1,0 this gives us a width of 2 at the focus

OpenStudy (amistre64):

so.... f(x) = a(-1-x)(1-x), such that f(0) = -3 -3 = a(-1)(1), a=3 our parabola to move is therefore f(x) = 3(-1-x)(1-x) , but how to move it?

OpenStudy (amistre64):

we need to raise it by 3, and move it to the right by 4

OpenStudy (amistre64):

so the alterations are: f(x) = 3(-1-(x-4))(1-(x-4)) + 3 and from that we can simplify it to our hearts content

OpenStudy (ulahlynn):

ohh because of axis x=4?

OpenStudy (amistre64):

yes, and i have to review my setup since i evidently got something off along the way

OpenStudy (amistre64):